# ARCO:Adaptive Multi-Agent Reinforcement Learning-Based Hardware/Software Co-Optimization Compiler for Improved Performance in DNN Accelerator Design

### v2

@@ 3.1. Overview of ARCO
Figure  2 depicts the architecture of our sophisticated co-optimizing compiler, termed ARCO, which elucidates the process of optimization compilation. ARCO initiates with a code template, $\tau$, for each layer of the neural network and a corresponding design space $D_{\Theta}$, delineating the range of possible configurations. The compiler then engages in an optimization process, navigating through the configuration space $\Theta$ to ascertain the optimal code configuration, represented as $\tau(\Theta^{*})$.
@@
ARCO maneuvers through the design space, employing a cost model as a surrogate for direct hardware measurements. The MARL Exploration module generates an initial set of candidate configurations $S_{\Theta}$. Through the CS method, ARCO refines $S_{\Theta}$ to produce a more focused subset $S^{\prime}_{\Theta}$, which contains a condensed yet highly promising set of configurations. The configurations in $S^{\prime}_{\Theta}$ are then passed to the MARL Code Generator module, which incorporates them with the input template $\tau$ to create a series of potential executable codes $\tau(\Theta)$. These are subsequently deployed on the hardware for empirical runtime evaluations. The hardware runtimes provide a measure of the configurations’ fitness, captured by a fitness function $f$, which informs the update of the cost model and enhances the exploration in subsequent iterations. After several iterations, the process converges to some $\tau(\Theta^{*})$ that achieves optimal fitness, characterized by the shortest hardware execution runtime. This configuration $(\Theta^{*})$ is then selected as the output for that network layer.
@@ 3.2. MARL Exploration
The MARL Exploration strategy employs three agents, each equipped with a policy network, as detailed in Table 1. Additionally, there is a centralized critic network (value network) that, along with the policy networks, is implemented as a Multi-Layer Perceptron (MLP) neural network. The policy network directs each agent to propose adjustments to the configuration knobs (as shown in Table 2) within its assigned portion of the design space. In contrast, the value network estimates the potential value of these adjustments. A shared central critic aids in evaluating the global state, facilitating collective learning during the training phase, while each agent makes decisions independently during execution, adhering to the CTDE paradigm.
@@
Software Agent 1
(Scheduling Agent)
This agent focuses on how tasks are parallelized and distributed, which is crucial for effective scheduling. Scheduling tasks typically involve deciding how computations are assigned and executed in parallel threads.
Software Agent 2
(Mapping Agent)
This agent deals with how data (namely, tensor height and width) is divided and processed. The focus on specific data dimensions makes this agent suitable for determining how computations map onto the hardware, dealing with spatial data distribution.
Hardware Agent This agent adjusts parameters that directly influence how different components of the tensors (batches, input channels, and output channels) are divided and processed by the hardware. This suggests a deep focus on optimizing the hardware utilization based on the hardware’s capabilities and limitations (namely, resource counts and performance levels.)
The optimization unfolds through several episodes, each consisting of multiple search steps. During these steps, the agents evaluate the current configuration and independently decide on actions to improve subsequent configuration fitness, guided by shared insights and individual observations. The process detailed in Algorithm 1 outlines how each agent interacts within the CTDE framework, iterating through configurations to enhance the overall network performance.
@@ 3.2.1. Cost Model and Central Critic Update
In the ARCO framework, the optimization goal is to improve throughput, with the cost model reflecting the inverse of execution time. The central critic, a key component of our MARL setup, updates its policy based on the aggregated feedback from all agents, continuously optimizing the global state evaluation. The critic’s learning process is guided by the update rule, which is mathematically expressed in section 2.2, which minimizes the mean squared error between the predicted values and the actual rewards obtained.
@@ 3.2.2. Incorporating Hardware and Software Constraints
This modification ensures that the MARL agents not only seek to optimize performance but also adhere to specified design constraints, balancing efficiency with practical deployment considerations.
@@
Hardware Design Agent (Hardware Optimizer) tile_b, tile_ci, tile_co
@@
Scheduling Agent (Software Optimizer) h_threading, oc_threading
@@
Mapping Agent (Software Optimizer) tile_h, tile_w
@@ 3.3. Confidence Sampling
In the domain of DNN optimization, the ability to efficiently sample configuration spaces without costly hardware measurements is essential for rapid compilation. We introduce the Confidence Sampling (CS) method. This technique is designed to leverage the probability distribution of configurations’ fitness judiciously to select a subset likely to yield high-performing network configurations.
Although the Confidence Sampling method operates on the principles of probability-guided selection and shares some conceptual similarities with Importance Sampling (IS) (Tokdar and Kass, 2010) in statistical methods, it is distinct in its purpose and application. In IS, samples are re-weighted based on their probability density to approximate expected values more efficiently. Confidence Sampling, however, specifically focuses on reducing the number of hardware measurements required when exploring the design space. By evaluating the configurations’ values, this method synthesizes a subset that not only encapsulates the diversity of the configuration space but also emphasizes the high-confidence regions, i.e., areas where the likelihood of encountering an optimal or near-optimal configuration is high.
@@
As part of the larger system, the CS method interacts closely with the exploration and exploitation mechanisms to ensure that the exploration is directed toward the most promising regions of the design space. By applying the CS method, we have observed a reduced number of required configurations, as shown in Figure  4. Moreover, we can observe the sampling process gravitates towards configurations that demonstrate superior performance over time, highlighting the efficacy of this method.
@@ 4.1. Experimental Setup
We have incorporated ARCO within the TVM framework to thoroughly evaluate its components, contrasting its performance with that of AutoTVM and CHAMELEON. This integration allows for a comprehensive, end-to-end assessment of ARCO across a suite of advanced deep-learning models, including AlexNet, VGG-11, VGG-13, VGG-16, VGG-19, ResNet-18, and ResNet-34. All models were benchmarked using the same AutoTVM compilation duration to ensure a rigorous and equitable comparison. Further, we have employed the source code of the CHAMELEON (Ahn et al., 2020) to compare its efficacy with AutoTVM and ARCO. As mentioned above, AutoTVM and CHAMELEON do not support hardware configuration exploration; thus, in the case of these two exploration techniques, we have considered the default specification values of the VTA++ (Banerjee et al., 2021). Performance tests were conducted on a robust platform powered by a 3.4 GHz AMD EPYC 7763 64-Core Processor, providing a stable and powerful environment for these evaluations. As discussed before, we used the VTA++ simulator as our target hardware.
Table 3 shows the details of the DNN models employed for the comparison studies. In our setup, each agent is equipped with its own policy network, and a centralized value network functions as the critic for all agents. Both networks are implemented as Multi-Layer Perceptrons (MLPs) using TensorFlow (Abadi et al., 2016) that their details are:
@@
Table 3. Details of the DNN models used in evaluating ARCO (All of the models are extracted from Mxnet (Chen et al., 2015)).
Network Dataset Number of Convolution Tasks
AlexNet(Krizhevsky et al., 2012) ImageNet 5
VGG-11(Simonyan and Zisserman, 2014) ImageNet 8
VGG-13(Simonyan and Zisserman, 2014) ImageNet 10
VGG-16(Simonyan and Zisserman, 2014) ImageNet 13
VGG-19(Simonyan and Zisserman, 2014) ImageNet 16
ResNet-18(He et al., 2016) ImageNet 17
ResNet-34(He et al., 2016) ImageNet 33
@@
Table 4. Hyper-parameters (HPs) used in ARCO.
@@ 4.2. End-to-end Evaluation
Figure 5 Compares the achieved throughput of different frameworks over AutoTVM. As the results illustrate, the significant reduction in mean inference time (advancement in throughput) achieved by ARCO compared to AutoTVM and CHAMELEON for various DNN models: AlexNet, VGG-11, VGG-13, VGG-16, VGG-19, and ResNet-18, and ResNet-34. On average, our methodology demonstrates a 1.17× improvement in throughput. Table 6 reports detailed numerical results supporting these findings.
@@
Model AutoTVM CHAMELEON ARCO
@@
Figure 6. Comparing the compilation time of different frameworks (The percentages show the speedup of ARCO compared to AutoTVM).
While achieving improvement in throughput, ARCO also needs less compilation time compared to AutoTVM; Figure 6 shows that ARCO speeds up the optimization time up to 42.2%. Figure 7 shows ARCO’s comparative output code performance relative to other frameworks. Notably, ARCO achieves convergence to the peak GFLOPS value, equivalent to those reached by AutoTVM and CHAMELEON (without adaptive sampling), but with greater efficiency. This is accomplished with fewer hardware measurements and at a faster rate, underscoring the efficacy of the CS method.
Figure 7. Comparing the compiled code performance of different frameworks for ResNet-18 model.
The overall efficiency gains are primarily attributable to the sophisticated MARL exploration techniques and the reduced hardware measurements necessitated by implementing the CS method.
@@ 5. Conclusion
This paper introduced ARCO, an innovative multi-agent reinforcement learning-based compiler that significantly advances the field of DNN accelerator design. By integrating a multi-agent system with specialized roles—two agents focused on software optimization and one on hardware—ARCO addressed the intricacies of DNN architecture and effectively navigated the vast configuration space to enhance performance and efficiency. Employing the MARL Exploration module and Confidence Sampling method, ARCO reduced the computational overhead and expedited the optimization process while increasing the throughput. Our experimental results, conducted on advanced DNN models like AlexNet, VGG, and ResNet series, demonstrated that ARCO outperforms traditional frameworks like AutoTVM and CHAMELEON, improving throughput by up to 37.95% and reducing optimization time significantly (Up to 42.2%). These enhancements underscore the potential of adaptive, intelligent systems in optimizing hardware and software layers cohesively, setting a new benchmark for future developments in real-time DNN applications.

### v3

@@ 3.1. Overview of DCOC
Figure  1 depicts the architecture of our sophisticated co-optimizing compiler, termed DCOC, which elucidates the process of optimization compilation. DCOC initiates with a code template, $\tau$, for each layer of the neural network and a corresponding design space $D_{\Theta}$, delineating the range of possible configurations. The compiler then engages in an optimization process, navigating through the configuration space $\Theta$ to ascertain the optimal code configuration, represented as $\tau(\Theta^{*})$.
@@
DCOC maneuvers through the design space, employing a cost model as a surrogate for direct hardware measurements. The MARL Exploration module generates an initial set of candidate configurations $S_{\Theta}$. Through the CS method, DCOC refines $S_{\Theta}$ to produce a more focused subset $S^{\prime}_{\Theta}$, which contains a condensed yet highly promising set of configurations. The configurations in $S^{\prime}_{\Theta}$ are then passed to the MARL Code Generator module, which incorporates them with the input template $\tau$ to create a series of potential executable codes $\tau(\Theta)$. These are subsequently deployed on the hardware for empirical runtime evaluations. The hardware runtimes provide a measure of the configurations’ fitness, captured by a fitness function $f$, which informs the update of the cost model and enhances the exploration in subsequent iterations. After several iterations, the process converges to some $\tau(\Theta^{*})$ that achieves optimal fitness, characterized by the shortest hardware execution runtime. This configuration $(\Theta^{*})$ is then selected as the output for that network layer.
@@ 3.2. MARL Exploration
The MARL Exploration strategy employs three agents, each equipped with a policy network, as detailed in Table 1. Additionally, there is a centralized critic network (value network) that, along with the policy networks, is implemented as a Multi-Layer Perceptron (MLP). The policy network directs each agent to propose adjustments to the configuration knobs (as shown in Table 2) within its assigned portion of the design space. In contrast, the value network estimates the potential value of these adjustments. A shared central critic aids in evaluating the global state, facilitating collective learning during the training phase, while each agent makes decisions independently during execution, adhering to the CTDE paradigm.
@@
Scheduling
Agent
Focuses on task parallelization and distribution for effective scheduling.
Mapping
Agent
Divides and processes data dimensions (tensor height and width) for optimal hardware computation mapping.
Hardware
Agent
Adjusts parameters for dividing and processing tensor components (batches, input and output channels) to optimize hardware utilization based on resource counts and performance levels.
The optimization unfolds through several episodes, each consisting of multiple search steps. During these steps, the agents evaluate the current configuration and independently decide on actions to improve subsequent configuration fitness, guided by shared and individual observations. The process detailed in Algorithm 1 outlines how each agent interacts within the CTDE, iterating through configurations to enhance the network performance.
@@ 3.2.1. Cost Model and Central Critic Update
In the DCOC framework, the optimization goal is to improve throughput, with the cost model reflecting the inverse of execution time. The central critic, a key component of our MARL setup, updates its policy based on the aggregated feedback from all agents, continuously optimizing the global state evaluation. The critic’s learning process is guided by the update rule, which is mathematically expressed in section 2.1, which minimizes the mean squared error between the predicted values and the actual rewards obtained.
@@ 3.2.2. Incorporating Hardware and Software Constraints
This modification ensures that the MARL agents not only seek to optimize performance but also adhere to specified design constraints, balancing efficiency with practical deployment considerations. In this work, without loss of generality, we considered VTA++ (Banerjee et al., 2021) as the HW platform. The VTA++ architecture is comparable to other hardware platforms, such as GPUs and FPGAs, due to its configurable nature and similarity in processing elements (Banerjee et al., 2021). This similarity ensures that the optimization strategies and results obtained from VTA++ can be generalized and applied to these other hardware platforms, validating the broad applicability of our framework.
@@
Hardware Design Agent
(Hardware Optimizer)
@@
Scheduling Agent
(Software Optimizer)
@@
Mapping Agent
(Software Optimizer)
@@ 3.3. Confidence Sampling
In DNN optimization domain, the ability to efficiently sample configuration spaces without costly hardware measurements is essential for rapid compilation. We introduce the Confidence Sampling (CS) method. This technique is designed to leverage the probability distribution of configurations’ fitness judiciously to select a subset likely to yield high-performing network configurations.
Although the Confidence Sampling method operates on the principles of probability-guided selection and shares some conceptual similarities with Importance Sampling (IS) (Tokdar and Kass, 2010) in statistical methods, it is distinct in its purpose and application. In IS, samples are re-weighted based on their probability density to approximate expected values more efficiently. Confidence Sampling, however, specifically focuses on reducing the number of hardware measurements required when exploring the design space. By evaluating the configurations’ values, this method synthesizes a subset that not only encapsulates the diversity of the configuration space but also emphasizes the high-confidence regions, i.e., areas where the likelihood of encountering an (near-) optimal configuration is high.
@@
By applying the CS method, we have observed a reduced number of required configurations, as shown in Figure  3. Moreover, we can observe the sampling process gravitates towards configurations that demonstrate superior performance over time, highlighting the efficacy of this method.
@@ 4.1. Experimental Setup
We have incorporated DCOC within the TVM framework to thoroughly evaluate its components, contrasting its performance with that of AutoTVM and CHAMELEON. This integration allows for a comprehensive, end-to-end assessment of DCOC across a suite of advanced deep-learning models, including AlexNet, VGG-11, VGG-13, VGG-16, VGG-19, ResNet-18, and ResNet-34. All models were benchmarked using the same AutoTVM compilation duration to ensure a rigorous and equitable comparison. Further, we have employed the source code of the CHAMELEON (Ahn et al., 2020) to compare its efficacy with AutoTVM and DCOC. As mentioned above, AutoTVM and CHAMELEON do not support hardware configuration exploration. Hence, we also added the hardware architecture agent for them to have a fair comparison. Performance tests were conducted on a robust platform powered by a 3.4 GHz AMD EPYC 7763 64-Core Processor, providing a stable and powerful environment for these evaluations. As discussed before, we used the VTA++ simulator as our target hardware.
In our setup, each agent is equipped with its own policy network, and a centralized value network is the critic for all agents. Both networks are implemented as Multi-Layer Perceptrons (MLPs) using TensorFlow (Abadi et al., 2016) that their details are:
@@
Table 3. Hyper-parameters (HPs) used in DCOC.
@@ 4.2. End-to-end Evaluation
Figure 4 Compares the achieved throughput of different frameworks over AutoTVM. As the results illustrate, the significant reduction in mean inference time (advancement in throughput) achieved by DCOC compared to AutoTVM and CHAMELEON for various DNN models: AlexNet, VGG-11, VGG-13, VGG-16, VGG-19, and ResNet-18, and ResNet-34. On average, our methodology demonstrates a 1.17× improvement in throughput. Table 5 reports detailed numerical results supporting these findings.
@@
Model AutoTVM CHAMELEON DCOC
@@
Figure 5. Comparing the compilation time of different frameworks (The percentages show the speedup of DCOC compared to AutoTVM).
While achieving improvement in throughput, DCOC also needs less compilation time compared to AutoTVM; Figure 5 shows that DCOC speeds up the optimization time (the entire process of compiling) up to 42.2%. Figure 6 shows DCOC’s comparative output code performance relative to other frameworks. Notably, DCOC achieves convergence to the peak GFLOPS value, equivalent to those reached by AutoTVM and CHAMELEON (without adaptive sampling), but with greater efficiency. This is accomplished with fewer hardware measurements and at a faster rate, underscoring the efficacy of the CS method.
Figure 6. Comparing the compiled code performance (GFLOPS per Trial) of different frameworks for ResNet-18 model.
@@ 5. Conclusion
This paper introduced DCOC, an innovative multi-agent reinforcement learning-based compiler that significantly advances the field of DNN accelerator design. By integrating a multi-agent system with specialized roles—two agents focused on software optimization and one on hardware—DCOC addressed the intricacies of DNN architecture and effectively navigated the vast configuration space to enhance performance and efficiency. Employing the MARL Exploration module and Confidence Sampling method, DCOC reduced the computational overhead and expedited the optimization process while increasing the throughput.